A bus leaves a station at 1PM, traveling west at an average rate of 44 mile per hour. One hour later a second bus leaves the same station, traveling east at a rate of 48 mile per hour. At what time will the two buses be 274 miles apart?
What formula should I use to find the exact answer? Just by manually adding it up I got the two would be at that distance around 4:32, but I don't know how to get to that answer with a formula.

Answer by ankor@dixie-net.com(12700)

A bus leaves a station at 1PM, traveling west at an average rate of 44 mile per hour. One hour later a second bus leaves the same station, traveling east at a rate of 48 mile per hour. At what time will the two buses be 274 miles apart?
:
Let t = travel time of the westbound bus, which left at 1pm
Then
(t-1) = travel time of the eastbound bus, which left at 2pm
:
Write a distance equation: Dist = speed * time
:
W.Bus dist + E.Bus dist = 274
44t + 48(t-1) = 274
:
44t + 48t - 48 = 274
:
92t = 274 + 48
:
t = 322/92
:
t = 3.5 hrs + 1 pm = 4:30 PM when they are 274 mi apart
:
:
Check solution:
44(3.5) + 48(2.5) = 274